OpenStudy (anonymous):

Solve three-sevenths plus five-ninths. Then use estimation to check for reasonableness. Show your work.

OpenStudy (texaschic101):

do you know how to add fractions ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

so what did you get for 3/7 + 5/9 ??

OpenStudy (anonymous):

8/16

OpenStudy (anonymous):

1/7 is smaller than 1/6. 5/6 is close of a whole number, but it is missing 1/6. Thus if we add 1/7 to 5/6, we should get a number that is ALMOST 1. ie. 42/42 Therefore by estimation, 41/42 is a correct answer.

OpenStudy (texaschic101):

3/7 + 5/9 --- common denominator is 63 27/63 + 35/63 = 62/63 estimation : 3/7 = 1/2 5/9 = 1/2 1/2 + 1/2 = 1 62/63 is very close to 1

OpenStudy (anonymous):

so thats the answer

OpenStudy (texaschic101):

well...when you add 3/7 + 5/9 you get 62/63.....almost 1 and when you estimate : 3/7 = 1/2 5/9 = 1/2 added together = 1 I think that is how you would write the answer

OpenStudy (anonymous):

or nah

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

sure thing :)